OpenStudy (anonymous):

Which fraction is equivalent to 5/ 6 + 7/ 8 ? A 35/ 48 B 6/ 7 C 20/ 21 D 41/ 24

OpenStudy (anonymous):

41/24 5/6 + 7/8 = (20+21)/24 = 41/24

OpenStudy (anonymous):

first find a common denominator

OpenStudy (anonymous):

thax

OpenStudy (anonymous):

then u are able to add

OpenStudy (anonymous):

so what is the LCD

OpenStudy (anonymous):

?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

24 is the LCD

OpenStudy (anonymous):

look

OpenStudy (anonymous):